OpenStudy (anonymous):

The graph below shows a line segment AB. Graph Which of the following equations best represents the line segment AB? y = -3x + 2 y = -2x + 1 y = -3x - 2 y = -2x - 1

OpenStudy (anonymous):

thinking A or C!

OpenStudy (anonymous):

@superdavesuper can you PPPPPPPPPPLLLLLLLLLLEEEEEEEAAAAAAAAAAAAASSSSSSSSSSSSSSEEEEEEEEEEE help?

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

why u think its A or C?

OpenStudy (anonymous):

they just look right

OpenStudy (superdavesuper):

No, i wont give u the ans anymore but i will help...if u start guessing the ans, i will leave. im serious -.-

OpenStudy (anonymous):

okay

OpenStudy (superdavesuper):

OK how are A and C different?

OpenStudy (anonymous):

(I guess a lot)

OpenStudy (superdavesuper):

Like I told u before, guessing do u very lil good :(

OpenStudy (anonymous):

A is -3x + 2 and C is -3x -2

OpenStudy (anonymous):

and I dont ask for the answers I just ask for the simplest way for yall to help me find it.

OpenStudy (superdavesuper):

no, u didnt ask for ans so u are good on that :) but u dont try to understand when i explain things to me n just start guessing n it makes me worry :(

OpenStudy (superdavesuper):

ok now back to A and C....one said +2 and one said -2.....how can we find out which one is right?

OpenStudy (anonymous):

Ummmmmmmmm...............plug in numbers????

OpenStudy (superdavesuper):

Good idea!! hehehe first time i see u trying here :) OK lets see what to plug in....do u see the vertical line in the middle where all the numbers 1,2,3,4,5 go?

OpenStudy (anonymous):

:P and yeps

OpenStudy (superdavesuper):

at that vertical line, x=0 okay?

OpenStudy (anonymous):

okay.

OpenStudy (superdavesuper):

So lets plug x=0 into A is -3x + 2 and C is -3x -2 and see what happens?

OpenStudy (anonymous):

okay so x is what? or is x were y goes bc in A and C it says y= -3x +2 and y= -3x +2

OpenStudy (anonymous):

finding x? are we?

OpenStudy (superdavesuper):

x=0

OpenStudy (superdavesuper):

x=0 then -3x + 2=?

OpenStudy (anonymous):

ooooooooooohhhhhhhhhhhh!!!!!!!!!!!!!!!! -1?

OpenStudy (superdavesuper):

no, x=0 then -3x + 2= -3(0) + 2 =?

OpenStudy (anonymous):

or with - number does it go more - ?

OpenStudy (anonymous):

WWWWAAAAAAAAAAIIIIIIIIIITTTTTTTTTT!!!!!! -3 time 0?

OpenStudy (superdavesuper):

it is a 0 inside and everything that multiplies 0 will be 0 right?

OpenStudy (superdavesuper):

yes, -3 times 0 indeed

OpenStudy (anonymous):

and than 0 + 2?

OpenStudy (superdavesuper):

U not reading my steps again :(

OpenStudy (superdavesuper):

yes so x=0 then -3x + 2=?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

BECAUSE -3 times 0 is 0 and 0 + 2 it 2 :)

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

so u now get x=0, y=2 right?

OpenStudy (anonymous):

yes!!!!!

OpenStudy (anonymous):

now what?

OpenStudy (superdavesuper):

READ this carefully!!!! (x,y) means a POINT on the graph u go to horizontal line in the middle and find x; then go vertical to y that point is ur (x,y) i cannot go on if u dont understand this

OpenStudy (anonymous):

okay, and on this graph I think x is a and y is b so I got it, I think. I kinda need to hurry because I have to leave in about 1 hour and I still have 4 questions left (after this)

OpenStudy (superdavesuper):

what is the right ans for this?

OpenStudy (anonymous):

(dont get mad if this isnt right) A?

OpenStudy (superdavesuper):

A is the right ans :)

OpenStudy (anonymous):

YYYYYEEEEAAAAHHHH!!!!!

OpenStudy (superdavesuper):

I am not mad...i AM worry about u thou. sorry I need to leave so good luck w the rest of ur assignments :)

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -3, 6, -9, -12, … -3, 20, -95, 480, … -3, 6, -9, 12, … -3, -20, -95, -480, …

OpenStudy (anonymous):

start another question?

OpenStudy (superdavesuper):

Please post it as a new one separately; hopefully someone else will help u. good luck!

OpenStudy (anonymous):

thanks :)

OpenStudy (superdavesuper):

welcome :)